Joe delivers 5 newspapers in the same time Matt delivers 4 newspapers. If they have a total of 54 papers, how many did Joe deliver? How many did Matt deliver?
Answer by ankor@dixie-net.com(22740)

Joe delivers 5 newspapers in the same time Matt delivers 4 newspapers.
If they have a total of 54 papers, how many did Joe deliver?
How many did Matt deliver?
:
let x = the multiplier
then
5x = number of papers delivered by Joe
and
4x = no. of papers by Matt
:
5x + 4x = 54
9x = 54
x = 54/9
x = 6 is the multiplier
therefore
6*5 = 30 delivered by Joe
and
6*4 = 24 delivered by Matt